# A Hierarchy of Spectral Gap Certificates for Frustration-Free Spin Systems

### 06 Nov 2024 v1

@@ Appendix E Certifiable formulation of the LTI SDP
s.t. $\quad g_{n}(\delta)+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}\succeq 0.$ (134)
$\quad g_{n,i}(\delta):=\left(h_{i}^{2}+\sum\limits_{i<j<i+n-k}\{h_{i},h_{j}\}-\delta h_{i}\right),$ (135)
the solver produces a solution $\delta$ for which the constraint is positive semidefinite only to some finite precision:
@@
for some small $\epsilon$ (for MOSEK, which is the solver we use, $\epsilon\approx 1e-8$). Such a solution cannot be used to prove a gap in the thermodynamic limit (or for large finite systems) because for a slightly negative $q_{n}$ as in eq. 136, we get
$\sum^{N}_{i}\tau_{i}q_{n}\succeq-N\epsilon$ (137)
. Therefore, to produce rigorous certificates for out of finite-precision solutions, we have to ensure that $q_{n}(\delta)$ is strictly positive semidefinite.
We can make such a certificate if the ground state of $H$ is given to us as a matrix-product-state. Note that for a frustration-free Hamiltonian the ground state is always an MPS if the ground state degeneracy remains bounded for all $N$ (theorem IV.17 in [48]). Moreover, in all practical use-case scenarios we have an MPS description of the ground state. We therefore assume that the MPS tensor describing the ground state is given to us.
To make our solution certifiable we slightly modify the LTI SDP. We proceed as follows: First we note that all the terms in $g_{n}$ annihilate the ground state MPS as the latter is annihilated by each individual term $h_{i}$. Note that, for $q_{n}$ to be a valid generating term for $[H^{2}]_{n}-\delta H$, the extra term $\mathbb{I}\otimes Y-Y\otimes\mathbb{I}$ also has to annihilate the ground state MPS. To ensure this, we construct the projector to the orthogonal complement of the $n-1$ site MPS, which we denote as $P_{n-1}^{\perp}$ (which is also the $(n-1)$-body parent Hamiltonian projector) and replace $Y$ with $P_{n-1}^{\perp}YP_{n-1}^{\perp}$ in eq. 133. This ensures that the whole operator $q_{n}$ annihilates the ground state MPS. The resulting $\delta$ might now be lower because we restricted the set of allowed $Y$s, but this is acceptable, since it still represents a lower bound on the gap. The new form for the certification LTI SDP is then
@@
With the above constraint, we have ensured that the new operator $q_{n}(\delta)$ annihilates the ground state MPS. This means, by construction $q_{n}(\mathbb{I}-P_{n}^{\perp})=0$ as the support of the projector $(\mathbb{I}-P_{n}^{\perp})$ is spanned by the $n$-body MPS (with all possible boundary states). Additionally, to fully ensure that $q_{n}(\delta)\succeq 0$ after the optimization is performed, we check the minimal eigenvalue of $\lambda_{\mathrm{min}}(P_{n}^{\perp}q_{n}(\delta)P_{n}^{\perp})$, which is $q_{n}$ projected to the orthogonal complement of the $n$-body MPS. If $\lambda_{\mathrm{min}}(P_{n}^{\perp}q_{n}(\delta)P_{n}^{\perp})$ is positive, then we have rigorously obtained a lower bound because the $q_{n}$ is zero in on $\mathbb{I}-P_{n}^{\perp}$ and strictly positive on $P_{n}^{\perp}$. On the other hand, if its smallest eigenvalue on $P_{n}^{\perp}$ is negative (we expect it to be small in magnitude because the constraint $q_{n}(\delta)$ can only be violated up to the solver precision) we keep reducing the gap ($\delta$) by a small number, until it results in a strictly positive smallest eigenvalue. If this is the case, we simply take the smaller value as our certified gap. In practice it is usually sufficient to introduce a correction on the order of $\epsilon/\gamma_{2}(h_{i})$, where $\gamma_{2}(h_{i})$ is the gap of the interaction term, to get a strictly positive $q_{n}$.

### 31 Mar 2026 v2

@@ Appendix D Certifiable formulation of the LTI SDP
s.t. $\quad g_{n}(\delta)+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}\succeq 0.$
$\quad g_{n}(\delta):=\left(h_{1}^{2}+\sum\limits_{1<j<1+n-k}\{h_{1},h_{j}\}-\delta h_{1}\right).$
We first demonstrate the problem of regarding a finite-precision solution as a proof. Typically the solver produces a solution $(\delta,Y)$ for which the constraint in eq. 112 is positive semidefinite only up to some finite precision $\epsilon>0$ :
@@
(for MOSEK, which is the solver we use, $\epsilon\approx 1e-8$). Such a solution cannot be used to prove a gap in the thermodynamic limit (or for large finite systems) because from $q_{n}$ as in eq. 113 we can only infer
$\sum^{N}_{i}q_{n,i}\succeq-N\epsilon.$ (114)
Therefore, to produce rigorous certificates out of finite-precision solutions, we need to find a way to deal with the solver’s inaccuracy.
To make our solution certifiable we slightly modify the LTI SDP: First we note that all the terms in $g_{n}$ annihilate any ground state of the $n$-body chain with open boundary conditions, as every ground state is annihilated by each individual term $h_{i}$. We choose to make the $Y$ terms in $q_{n}$ such that they also annihilate every ground state. To ensure this, we construct the projector to the orthogonal complement of the ground space of the n1n-1 site chain, which we denote as $P_{n-1}^{\perp}$, and replace $Y$ with $P_{n-1}^{\perp}YP_{n-1}^{\perp}$ in eq. 112. This ensures that the whole operator $q_{n}$ annihilates any ground state of the open chain. The optimal value $\delta$ found by the solver might now be lower because we restricted the set of allowed $Y$s, but this is acceptable, since it still represents a lower bound on the gap. The projectors $P_{n}$ can be either computed by exact diagonalization of the $n$-site Hamiltonian (we are always in the regime when this is possible as it is a subroutine needed to solve the SDP), or be given exactly by the matrix product state (MPS) representation of the ground state: the MPS with open bond indices spans the ground space of the open chain. The latter is the case for all the models we treated in this paper. The new form for the certifiable LTI SDP is then
@@
With the above modification we have ensured that, independently of the value of $\delta$, the kernel of $q_{n}(\delta)$ contains all the ground states of the open chain Hamiltonian. Other zero (or small negative) eigenvalues of the optimal $q_{n}(\Delta^{\textbf{cert}}_{\mathrm{LTI}}(n))$ found by the solver arise from the maximization over $\delta$ as part of the SDP solution. We can then check the minimal eigenvalue of $q_{n}$ outside of its kernel and try to make it positive by modifying the gap estimate $\Delta^{\textbf{cert}}_{\mathrm{LTI}}(n)$ produced by the solver. For example, we can perform a line search over $\eta\in(0,1]$, halting the search when $\lambda_{\mathrm{min}}(P_{n}^{\perp}q_{n}(\Delta^{\textbf{cert}}_{\mathrm{LTI}}(n)-\eta)P_{n}^{\perp})$ is bigger than some positive threshold. In practice it is usually sufficient to introduce a correction on the order of $\epsilon/\gamma_{2}(h_{i})$, where $\gamma_{2}(h_{i})$ is the gap of the interaction term, to get a strictly positive $q_{n}$.